# Impact of evanescent scattering modes and finite dispersion in superconducting junctions

### v1

@@ IV Beenakker equation including evanescent modes
Now, if dispersion is negligible, i.e. $\left|\mu_{n}\right|\gg\Delta,E$, and considering that by definition evanescent (planar) modes have $\mu_{n}>0$ ($\mu_{n}<0$), Eqs. (66) and (67) imply that for evanescent momenta ${K_{\mathrm{e}}\approx K_{\mathrm{h}}\approx k_{\mathrm{I}}}$ and ${k_{\mathrm{R}}\approx 0}$, whereas for planar momenta ${k_{\mathrm{e}}\approx k_{\mathrm{h}}\approx k_{\mathrm{R}}}$ and ${k_{\mathrm{I}}\approx 0}$. Going to the limit of small $\Delta,E$ with respect to $\mu_{n}$, the explicit expressions (62) and (63) can be expanded up to first order as follows. For planar channels $n\leq n_{\text{P}}$ (i.e., for the respective “pp” subblocks) we get the entries
@@
Indeed, in the limit of $\Delta/\mu_{n}\rightarrow 0$, we obtain the standard framework (Sec. II), where both dispersive effects and evanescent modes are irrelevant, as here $\widehat{\alpha}_{\nu}^{\text{pp}}=\alpha$, while $\widehat{\alpha}_{\nu}^{\text{ee}}=0$ as well as $\widehat{\beta}_{\nu}=0$. Overall, we clearly see the very distinct behavior between the planar and evanescent coefficients. For the planar coefficients, both regular Andreev reflection and normal backscattering have a first order correction. For the evanescent coefficient, there only is a weak, nonzero Andreev reflection, whereas normal backscattering is absent. Moreover, both planar coefficients explicitly depend on energy $E$ (note that Eqs. (68) and (69) contain the energy-dependent factor $\alpha$), whereas evanescent Andreev reflection is constant.
@@
Crucially, the above expansion also establishes a concrete guideline for assessing which modes can be neglected and which ones must be included. For a dispersion relation as shown in Fig. 4 there is only one mode with a small $\mu_{n}$, namely the one whose energy minimum is closest to the chemical potential $\mu$. When tuning the chemical potential just below the channel number transition, the total number of modes that need to be included is $n_{\text{P}}+1$, i.e., all planar modes, plus the lowest evanescent mode with strong dispersion. With a chemical potential above the transition, we again only have to take into account the planar modes ($n_{\text{P}}$), but include finite dispersive effects of the highest planar mode (in analogy to Sec. III.3). This very narrow cut-off is insofar surprising as one might have expected that every evanescent mode with a sufficiently long exponential tail (such that it penetrates the SN-interface) would have to be included. For extremely short junctions, this would have concerned a high number of modes. The above expansion result, on the other hand, shows that penetration of the SN-interface only is a necessary condition. In addition, modes also need to have finite dispersion, i.e. ${\left|\mu_{n}\right|\approx\Delta,E}$, in order to be relevant. If not, backscattering is negligible and the planar reflection coefficient $\alpha_{\nu,n}$ reduces to the conventional form, Eq. (12) (whereas evanescent Andreev reflection vanishes completely), such that the interplay of energy dependent scattering and evanescent Andreev reflection, as discussed above, does not apply. As argued above, in situations where only one mode is close to a transition point, all modes higher than the first evanescent channel usually can be neglected. This is a further central result of our work. We will illustrate this principle, and its implications, explicitly in the subsequent sections.
@@
To summarize, within the new framework, Andreev reflection and scattering are still governed by the boundary conditions as expressed in Equations (1) and (10), such that Eq. (17) remains valid. However, the structure of both $S$ and $R$ is changed (both including planar as well as evanescent modes). Consequently, when expressing the interference condition in terms of the submatrix structure, we get an equation of much more general form than, e.g., Eq. (18). In particular, by including backscattering, the reflection matrix $R$ loses its block off-diagonal form, as indicated by Eq. (57). Moreover, in contrast to the state of the art model, the reflection and backscattering coefficients $\widetilde{\alpha}_{\nu,n}$ and $\widetilde{\beta}_{\nu,n}$ are now different for each channel and, thus, can no longer be factored out. Consequently, Eq. (18) takes on the generalized form
@@ V Scattering matrix constraint including evanescent modes
whose size is given by the number of planar modes. While this matrix still fulfills particle-hole symmetry, $\tau_{y}\left[S_{\text{eff}}^{pp}(-E)\right]^{*}\tau_{y}=S_{\text{eff}}^{pp}(E)$ (where $\tau_{y}$ is the Pauli matrix in Nambu space), probability conservation is not guaranteed, hence, it cannot be considered a true scattering matrix.
The truncation to planar levels in terms of this pseudo-scattering matrix is nonetheless very instructive to analyze the relationship between evanescent scattering and band touching, in a similar fashion as for purely planar processes in Sec. III.3. To that end, note that this effective scattering matrix is no longer block-diagonal in the electron-hole space. This is due to the fact that $R^{\text{ee}}$ has finite Andreev reflection, due to the finite dispersion of the evanescent mode, Eq. (70). It now turns out that the interplay between diagonal and off-diagonal processes (electron-electron versus electron-hole, etc.) allows us to derive a condition relating the energy dependence of the planar part of the scattering matrix ($S^{\text{pp}}$) to the planar-evanescent ($S^{\text{ep}}$) and evanescent-planar ($S^{\text{pe}}$) processes, which is in its nature, crucially, beyond probability conservation arguments.
@@
$h_{\text{eff}}\left(E\right)=-i\Delta\left(\begin{array}[]{cc}\frac{\Delta}{8\mu_{2}}\Sigma&-E(\partial S_{e}^{\text{pp}})^{*}\\ E\partial S_{e}^{\text{pp}}&-\frac{\Delta}{8\mu_{2}}\Sigma^{*}\end{array}\right)\ ,$ (77)
@@
where (in analogy to Sec. III.3) all scattering matrices are evaluated (expanded) around $E=0$. This Hamilonian shares a striking structural similarity with Eq. (45), except that the subblock $S_{e}-S_{e}^{*}$ (representing the deformation of an incoming planar Gaussian wave packet due to finite dispersion) is here replaced with $\Sigma/2$ (due to the coupling between evanescent and planar modes). Notice furthermore, that $\Sigma$ satisfies the symmetry
@@
$\partial S_{e}^{\text{pp}}(S_{e}^{\text{pp}})^{*}-S_{e}^{\text{pp}}(\partial S_{e}^{\text{pp}})^{*}+\frac{1}{4\mu_{2}}\Sigma^{*}=0\ ,$ (80)
is satisfied. In this condition, the superconducting gap $\Delta$ is likewise eliminated, such that this is a condition exclusively concerning the scattering matrix (including planar and evanescent modes), irrespective of the presence or absence of Andreev processes. This is a further central result of our work. While much of the above derivation works in complete analogy to Sec. III.3, the two conditions of Eqs. (43) and (80) are fundamentally different in nature for the following reason. In Sec. III.3, it was possible to relate the conditions for a gap closing to probability conservation of the scattering process itself. Here, this is explicitly impossible, as the correction in Eq. (74) stems from virtual excitations of the electrons into the evanescent states, requiring both a coupling to decaying and diverging modes (i.e., the “detour” via evanescent modes involves both $S^{\text{pe}}$ and $S^{\text{ep}}$), only the former of which are normalizable (as we already discussed in Sec. III.5). Hence, even though the condition concerns only the properties of the scattering matrix (the parameters of the Andreev reflection have been eliminated), they do not refer to a simple probability conservation principle. Therefore, as already foreshadowed in Sec. III.5, while it is straightforward to anticipate from qualitative reasoning that there should exist a gap closing due to simple dwelling time arguments, the actual demonstration of this fact is difficult. Here, we have managed to find such a nontrivial relationship for the scattering matrix including evanescent modes by means of Andreev physics.
Finally, we emphasize again that condition (80) is guaranteed to be satisfied as $|\mu_{2}|$ approaches zero. In this regime, the length of the lowest evanescent mode diverges, and thus becomes much larger than the size of the scattering region (which will also contribute to a finite energy dependence of the planar part of the scattering matrix). In the next section, we consider a concrete scattering problem, where we can numerically confirm the validity of Eq. (80). Moreover, we will be able to show that, strikingly, the gap closing also holds beyond the perturbative limit, i.e., when neither $\mu_{2}$ nor $\Delta$ are small.
@@ VI Geometric scattering and L-junction model
As announced, the above framework, as well as the nontrivial aspects regarding the detaching of the ABS spectrum at the superconducting gap $\Delta$ can be illustrated with a simple, yet experimentally highly relevant example device with non-trivial geometry of the scattering center. Specifically, we consider a 2D conductor in a two terminal configuration with horizontal and vertical arms forming an inverted L-profile (c.f. Fig. 7[43, 44, 45]. In the absence of an additional scalar potential (or impurities), here, the scattering is completely dominated by the geometry of the device (henceforth referred to as geometric scattering), as it mixes the longitudinal and transversal components of the wave function at the scattering region. This makes it an ideal and well-controllable example for a scattering scenario where translational invariance is absent. While we here focus on a two-terminal device for simplicity, note that for generic multiterminal junctions, geometric scattering can always be expected to be present.
Figure 7: Discrete lattice model of a two-terminal junction where scattering occurs across an inverted L-geometry. Owing to the junction geometry, translational invariance is broken, such that there is significant cross-channel coupling. Lattice points that are proximitized by superconducting regions are marked in blue. These regions extend over ${j_{x},j_{y}\in\left[-\infty,j_{\alpha}\right]}$, where $j_{\alpha}$ denotes the length of the conductor arms in between the scattering center (${j=0}$) and the SN-interface (${j=j_{\alpha}}$). Transversal wave function components are confined by the width of the conductor arms, where $N$ is the number of transversal lattice points. In order to compute the scattering matrix, it is useful to consider cross-shaped plaquettes linking neighbouring lattice points due to the tight-binding tunnel hopping.
VI.1 Normal metal scattering with evanescent modes

### v2

@@ IV Beenakker equation including evanescent modes
Now, if dispersion is negligible, i.e., $\left|\mu_{n}\right|\gg\Delta,E$, and considering that by definition evanescent (planar) modes have $\mu_{n}>0$ ($\mu_{n}<0$), Eqs. (52) and (53) imply that for evanescent momenta ${\kappa_{\mathrm{e}}\approx\kappa_{\mathrm{h}}\approx k_{\mathrm{I}}}$ and ${k_{\mathrm{R}}\approx 0}$, whereas for planar momenta ${k_{\mathrm{e}}\approx k_{\mathrm{h}}\approx k_{\mathrm{R}}}$ and ${k_{\mathrm{I}}\approx 0}$. Going to the limit of small $\Delta,E$ with respect to $\mu_{n}$, the explicit expressions (48) and (49) can be expanded up to first order as follows. For planar channels $n\leq n_{\text{P}}$ (i.e., for the respective “pp” subblocks) we get the entries
@@
Indeed, in the limit of $\Delta/\mu_{n}\rightarrow 0$, we obtain the standard framework (Sec. II), where both dispersive effects and evanescent modes are irrelevant, as here $\widehat{\alpha}_{\nu}^{\text{pp}}=\alpha$, while $\widehat{\alpha}_{\nu}^{\text{ee}}=0$ as well as $\widehat{\beta}_{\nu}=0$. Overall, we clearly see the very distinct behaviour between the planar and evanescent coefficients. For the planar coefficients, both regular Andreev reflection and normal backscattering have a first order correction. For the evanescent coefficient, there only is a weak, nonzero Andreev reflection, whereas normal backscattering is absent. Moreover, both planar coefficients explicitly depend on energy $E$ (note that Eqs. (54) and (55) contain the energy-dependant factor $\alpha$), whereas evanescent Andreev reflection is constant.
@@
Crucially, the above expansion also establishes a concrete guideline for assessing which modes can be neglected and which ones must be included. For a dispersion relation as shown in Fig. 5 there is only one mode with a small $\mu_{n}$, namely the one whose energy minimum is closest to the chemical potential $\mu$. When tuning the chemical potential just below the channel number transition, the total number of modes that need to be included is $n_{\text{P}}+1$, i.e., all planar modes, plus the lowest evanescent mode with strong dispersion. With a chemical potential above the transition, we again only have to take into account the planar modes ($n_{\text{P}}$), but include finite dispersive effects of the highest planar mode (in analogy to Sec. III.3). This very narrow cut-off is insofar surprising as one might have expected that every evanescent mode with a sufficiently long exponential tail (such that it penetrates the SN-interface) would have to be included. For extremely short junctions, this would have concerned a high number of modes. The above expansion result, on the other hand, shows that penetration of the SN-interface only is a necessary condition. In addition, modes also need to have finite dispersion, i.e. ${\left|\mu_{n}\right|\approx\Delta,E}$, in order to be relevant. If not, backscattering is negligible and the planar reflection coefficient $\alpha_{\nu,n}$ reduces to the conventional form, Eq. (12) (whereas evanescent Andreev reflection vanishes completely), such that the interplay of energy dependant scattering and evanescent Andreev reflection, as discussed above, does not apply. As argued above, in situations where only one mode is close to a transition point, all modes higher than the first evanescent channel usually can be neglected. This is a further central result of our work. We illustrate this principle, and its implications, explicitly in the subsequent sections.
@@
To summarise, within the new framework, Andreev reflection and scattering are still governed by the boundary conditions as expressed in Equations (1) and (10), such that Eq. (17) remains valid. However, the structure of both $S$ and $R$ is changed (both including planar as well as evanescent modes). Consequently, when expressing the interference condition in terms of the submatrix structure, we get an equation of much more general form than, e.g., Eq. (18). In particular, by including backscattering, the reflection matrix $R$ loses its block off-diagonal form, as indicated by Eq. (43). Moreover, in contrast to the state of the art model, the reflection and backscattering coefficients $\widetilde{\alpha}_{\nu,n}$ and $\widetilde{\beta}_{\nu,n}$ are now different for each channel and, thus, can no longer be factored out. Consequently, Eq. (18) takes on the generalised form
@@ V Scattering matrix constraint including evanescent modes
Both the S- and the R-matrix can be cast into a block form separating explicitly planar and evanescent modes,
@@
whose size is given by the number of planar modes. While this matrix still fulfils particle-hole symmetry, $\tau_{y}\left[S_{\text{eff}}^{pp}(-E)\right]^{*}\tau_{y}=S_{\text{eff}}^{pp}(E)$ (where $\tau_{y}$ is the Pauli matrix in Nambu space), probability conservation is not guaranteed, hence, it cannot be considered a true scattering matrix.
The truncation to planar levels in terms of this pseudo-scattering matrix is nonetheless very instructive to analyse the relationship between evanescent scattering and band touching, in a similar fashion as for purely planar processes in Sec. III.3. To that end, note that this effective scattering matrix is no longer block-diagonal in the electron-hole space. This is due to the fact that $R^{\text{ee}}$ has finite Andreev reflection, due to the finite dispersion of the evanescent mode, Eq. (56). It now turns out that the interplay between diagonal and off-diagonal processes (electron-electron versus electron-hole, etc.) allows us to derive a condition relating the energy dependence of the planar part of the scattering matrix ($S^{\text{pp}}$) to the planar-evanescent ($S^{\text{ep}}$) and evanescent-planar ($S^{\text{pe}}$) processes, which is in its nature, crucially, beyond probability conservation arguments.
@@
$h_{\text{eff}}\left(E\right)=-i\Delta\left(\begin{array}[]{cc}\frac{\Delta}{8\mu_{2}}\Sigma&-E(\delta S_{e}^{\text{pp}})^{*}\\ E\delta S_{e}^{\text{pp}}&-\frac{\Delta}{8\mu_{2}}\Sigma^{*}\end{array}\right)\ ,$ (63)
@@
where (in analogy to Sec. III.3) all scattering matrices are evaluated (expanded) around $E=0$. This Hamiltonian shares a striking structural similarity with Eq. (31), except that the subblock $S_{e}-S_{e}^{*}$ (representing the deformation of an incoming planar Gaussian wave packet due to finite dispersion) is here replaced with $\Sigma/2$ (due to the coupling between evanescent and planar modes). Notice furthermore, that $\Sigma$ satisfies the symmetry
@@
$\delta S_{e}^{\text{pp}}(S_{e}^{\text{pp}})^{*}-S_{e}^{\text{pp}}(\delta S_{e}^{\text{pp}})^{*}+\frac{1}{4\mu_{2}}\Sigma^{*}=0\ ,$ (66)
is satisfied. In this condition, the superconducting gap $\Delta$ is likewise eliminated, such that this is a condition exclusively concerning the scattering matrix (including planar and evanescent modes), irrespective of the presence or absence of Andreev processes. This is a further central result of our work. While much of the above derivation works in complete analogy to Sec. III.3, the two conditions of Eqs. (29) and (66) are fundamentally different in nature for the following reason. In Sec. III.3, it was possible to relate the conditions for a gap closing to probability conservation of the scattering process itself. Here, this is explicitly impossible, as the correction in Eq. (60) stems from virtual excitations of the electrons into the evanescent states, requiring both a coupling to decaying and diverging modes (i.e., the “detour” via evanescent modes involves both $S^{\text{pe}}$ and $S^{\text{ep}}$), only the former of which are normalisable (as we already discussed in Sec. III.5). Hence, even though the condition concerns only the properties of the scattering matrix (the parameters of the Andreev reflection have been eliminated), they do not refer to a simple probability conservation principle. Therefore, as already foreshadowed in Sec. III.5, while it is straightforward to anticipate from qualitative reasoning that there should exist a gap closing due to simple dwelling time arguments, the actual demonstration of this fact is difficult. Here, we have managed to find such a nontrivial relationship for the scattering matrix including evanescent modes by means of Andreev physics.
Finally, we emphasise again that condition (66) is guaranteed to be satisfied as $|\mu_{2}|$ approaches zero. In this regime, the length of the lowest evanescent mode diverges, and thus becomes much larger than the size of the scattering region (which will also contribute to a finite energy dependence of the planar part of the scattering matrix). In the next section, we consider a concrete scattering problem, where we can numerically confirm the validity of Eq. (66). Moreover, we will be able to show that, strikingly, the gap closing also holds beyond the perturbative limit, i.e., when neither $\mu_{2}$ nor $\Delta$ are small.
@@ VI Geometric scattering and L-junction model
Figure 7: Discrete lattice model of a two-terminal junction where scattering occurs across an inverted L-geometry. Owing to the junction geometry, translational invariance is broken, such that there is significant cross-channel coupling. Lattice points that are proximitised by superconducting regions are marked in blue. These regions extend over ${j_{x},j_{y}\in\left[-\infty,j_{\alpha}\right]}$, where $j_{\alpha}$ denotes the length of the conductor arms in between the scattering centre (${j=0}$) and the SN-interface (${j=j_{\alpha}}$). Transversal wave function components are confined by the width of the conductor arms, where $N$ is the number of transversal lattice points. In order to compute the scattering matrix, it is useful to consider cross-shaped plaquettes linking neighbouring lattice points due to the tight-binding tunnel hopping.
As announced, the above framework, as well as the nontrivial aspects regarding the detaching of the ABS spectrum at the superconducting gap $\Delta$ can be illustrated with a simple, yet experimentally highly relevant example device with non-trivial geometry of the scattering centre. Specifically, we consider a 2D conductor in a two terminal configuration with horizontal and vertical arms, such that scattering occurs at an inverted L-profile (cf. Fig. 7), which is strongly inspired by recent experimental setups [44, 45, 46]. In the absence of an additional scalar potential (or impurities), here, the scattering is completely dominated by the geometry of the device (henceforth referred to as geometric scattering), as it mixes the longitudinal and transversal components of the wave function at the scattering region. This makes it an ideal and well-controllable example for a scattering scenario where translational invariance is absent. While we here focus on a two-terminal device for simplicity, note that for generic multiterminal junctions, geometric scattering can always be expected to be present.
Figure 9: Square root behaviour of the planar-planar reflection coefficient as a function of energy. Here, energies are taken with respect to the reference energy scale $\Delta$VI.The chemical potential $\mu/\Delta=98.99$ is tuned such that the second mode is evanescent (${\mu_{2}=\mu_{\mathrm{crit}}>0}$) and has strong non-linear dispersion with $\mu_{2}/\Delta=1.01$. The planar-planar reflection coefficient is computed numerically (dashed curve) with $N=5$ transversal lattice points for different energies (the plot shows its absolute squared $\left|R_{1,1}\right|^{2}$). Although the planar channel has negligible dispersion ($\mu_{1}/\Delta=-72.20$), energy dependence is strong due to cross-channel scattering with the second mode and follows a square root law (solid curve), in accordance with Eq. (91).